### v1

# Adaptively Optimizing the Performance of HPX’s Parallel Algorithms

Abstract
C++ Executors simplify the development of parallel algorithms by abstracting concurrency management across hardware architectures. They are designed to facilitate portability and uniformity of user-facing interfaces; however, in some cases they may lead to performance inefficiencies due to suboptimal resource allocation for a particular workload or not leveraging certain hardware-specific capabilities.
To mitigate these inefficiencies, we have developed a strategy based on cores and chunking (workload), and integrated it into HPX’s executor API. This strategy dynamically optimizes for workload distribution and resource allocation based on runtime metrics and overheads. In this paper, we introduce the model behind this strategy and evaluate its efficiency by testing its implementation (as an HPX executor) on both compute-bound and memory-bound workloads. The results show speedups across all tests, configurations, and workloads studied, offering improved performance through a familiar and user-friendly C++ executor API.
Keywords:
Asynchronous Many-Task (AMT) HPX Executors Parallel Algorithms.
1 Introduction
Managing parallelism and concurrency has become increasingly complex in software due to the heterogeneity of modern architectures. HPX provides executors to offer uniform APIs for execution of tasks across different architectures. While this abstraction eases development, it delegates many runtime metrics, such as number of cores or chunk sizes, to the execution policy, where poor choices can lead to performance degradation. For instance, even a straightforward embarrassingly parallel operation on a multi-core processor can be sensitive to the resource allocation strategy. While distributing work across all cores may be effective for larger input data, utilizing more cores for smaller input sizes may lead to performance degradation.
Beyond that, there is a wide range of parallel algorithm types, including map-type algorithms (e.g. copy, fill, stencil updates, etc.), map-reduce-type algorithms (e.g. min_element, all_of, count, and prefix sums, etc.), to name a few. The performance of these algorithms can vary depending on several factors, such as the nature of the algorithm (whether it is compute-bound or memory-bound), the number of cores and threads, data partitioning, and the specific architecture (cache sizes, memory bandwidth, etc.). The challenge is that no single set of parameters can guarantee optimal performance across all scenarios. The optimal configuration for parallel execution is often highly dependent on the specific workload and runtime environment. In the context of parallelization, there are four factors to consider: Starvation:, Latencies:, Overheads: and, Waiting for contention resolution [1]
This paper addresses these challenges by proposing a dynamic optimization approach integrated into HPX’s executor API. We propose an aggregated model which combines all the effects such as architectural specifics to get the best possible performance regarding the number of cores and chunk sizes. Our focus is on optimizing workload distribution and resource allocation through runtime adaptivity. This method aims to reduce overheads, such as those associated with parallelism management, while maintaining the simplicity of user-friendly APIs. We investigate the impact of different core counts and chunk sizes on two algorithms: adjacent-difference and artificial-work.
We use HPX, a C++ standard library for parallelism and concurrency, as the framework for this study, because it is easily customizable and open-source. Furthermore, it offers similar semantics and performance to OpenMP scheduling for comparable problems. Indeed, HPX has even been used to build a backend for OpenMP [2].
Our results demonstrate that adapting execution configurations based on runtime metrics can prevent the performance degradation often associated with generic abstractions like executors or OpenMP’s parallel for pragma.
2 Motivation
Scientific simulation codes are often based on Cauchy problems, i.e., codes described by an initial data on an N-dimensional spatial grid, and accompanied by a system of differential equations used to iteratively evolve the data through time. These codes, common in science and industry, typically involve the use of finite differences to simulate spatial derivatives, operations that are very much like the adjacent-difference algorithm which calculates the difference between each pair of elements within a sequence. Such algorithms are inherently parallel and can benefit from parallelization of loops or regions of independent tasks. To parallelize loops, the computational grid is split into chunks (chunking), such that each chunk can be independently updated on a separate core. Determining the size of the chunks and a reasonable number of processing units to obtain the best possible execution times is—more often than not—left to the programmer. These metrics are application-specific and require measurement and tuning for each of the parallel constructs involved. Common off-the-shelf runtimes like OpenMP and the C++17 Standard library expect the user to tune and provide such metrics. The effectiveness of a parallelization strategy also depends on system and implementation-specific characteristics, e.g., the overhead per scheduled task, the amount of (possibly nested) parallelism intrinsic to the application, the underlying computing architecture, and many more factors [3].
To address this loop-tuning problem, we will develop a methodology for adaptively determining the optimal chunk sizes and the number of processing units and implement it for HPX’s parallel algorithms.
3 Theory
What is the ideal number of cores to use for a given workload?
To start, we define $T_{1}$ as the total time taken to execute a loop on a single thread without parallelism, and $T_{N}$ as the total time with $N$ number of threads (assuming $N>1$). We now assume that the loop can be perfectly parallelized apart from a constant overhead, $T_{0}$.
$T_{N}=\frac{T_{1}}{N}+T_{0}$ (1)
Alternatively, we can derive the relation between speedup (which measures how much faster a parallel algorithm runs in compare to a sequential one), efficiency (which measures how well a system utilizes parallelism), and number of processors to calculate the optimal number of processors from first principles.
We can now compute speedup as follows:
$S=\frac{T_{1}}{T_{N}}$ (2)
Thus far, the speedup formula is generic to any model of computation, including Amdahl’s Law[4] and Gustafson’s Law[5]. However, substituting Equation 1 into Equation 2, we obtain the following:
$S=\frac{T_{1}}{\frac{T_{1}}{N}+T_{0}}$ (3)
We point out that this “Overhead Law” is different than Amdahl’s Law and Gustafson’s Law . The former assumes that a fixed fraction of the code is serial, the latter assumes a fixed amount of serial code is always present. In our case, we assume that a fixed amount of serial code is only present when parallelism is attempted. If one wished to create a formula more comparable to these other two laws, one could write the parallel fraction for our case as $p=T_{1}/(T_{0}+T_{1})$. Substituting this yields:
$S=\frac{p}{1-p+\frac{p}{N}}$ (4)
So the "Overhead Law" is differs from Amdahl’s Law by a constant factor. We point out that, unlike the other laws, this equation is not valid for the case $N=1$ since Equation 1 only applies when $N>1$.
Now that we have the “Overhead Law,” we can compute the efficiency, which is defined as the ratio of speedup to the number of threads:
$E=\frac{S}{N}$ (5)
Based on Equation 2 and 5 we can derive the following:
$E=\frac{S}{N}=\frac{T_{1}}{NT_{N}}$ (6)
Using Equation 1 and substituting $T_{N}$ based on $T_{1}$, simplifying them, we can solve it for N:
$N=\frac{1-E}{E}*\frac{T_{1}}{T_{0}}$ (7)
We will choose an efficiency (E) of 95% as before.
Note that, using Eq. 7, we can determine that $T_{opt}=19T_{0}$ in Eq. 8. In other words, if we allocate chunks of size $19T_{o}$ to each core, we will achieve $95\%$ efficiency.
In summary, this approach allows us to calculate $N_{C}$, the number of cores that we can make use of without exceeding our parallel efficiency goal.
Once the optimal core count for a given workload is determined, we need to find the optimal chunk size. To find that, We apply the following formula to determine the ideal amount of work per core:
$N_{C}\>=\>{\frac{T_{1}}{T_{opt}}}$ (8)
where as before, $T_{1}$ is the total time taken per workload and $T_{opt}$ is the time for minimum work per core. If we find the amount of work where parallelization with 2 cores gives a speedup of $1.9$ (a $95\%$ efficiency), based on our experiments we can, with the given formula, calculate $T_{opt}$. This optimal time represents the amount of work per core for $95\%$ efficiency. we derive the following formula using Eq. 8 to calculate the minimum workload for each chunk:
$T_{m}\>=\>{\frac{T_{1}}{N_{C}*C}}$ (9)
Where $N_{C}$ is the optimal number of cores calculated in Eq. 8, $C$ is chunks-per-core (which is equal to 8 based on the experiments), $T_{1}$ is the overall time per workload and $T_{m}$ is the amount of work per chunk. Assuming that $T_{1}$ and $T_{m}$ are both proportional to the number of elements they contain, we can derive an equation for the chunk size ($N_{CH}$). The value of $N_{E}$ is the number of elements in our workload.
$N_{CH}\>=\>{\frac{N_{E}}{N_{C}*C}}$ (10)
This equation ensures that $C=8$ chunks per core are used for any workload, with the chunk size always being at least $T_{m}$. In section 4 we will show how this time parameter is utilized to design the chunk size.
4 Implementation
Now that we know how to compute the ideal number of cores, $N_{C}$, we turn to the implementation. We use HPX because it is open source, easily customizable, and well-optimized. For this work, we will consider the parallelization of algorithms and loops, scheduling work in a manner similar to the familiar static scheduling of OpenMP.
We will start this section with some background on HPX, executors, and customization points.
4.1 HPX Algorithms and Customization
HPX is a C++ Standard Library for parallelism and concurrency [6, 7]. HPX is implemented as a lightweight user-level task manager running on top of kernel threads [8].
Execution policies are objects that can be provided as the first argument to the standard algorithms of C++. The C++17 and C++20 Standards introduce four execution policy types [9, 10].
  • std::execution::seq requires that a parallel algorithm’s execution not be parallelized.
  • std::execution::par indicates that a parallel algorithm’s execution may be parallelized. Within this policy, all the different scheduling types
    from OpenMP are available through objects called execution parameters.
  • std::execution::unseq indicates that a parallel algorithm’s execution may be vectorized, e.g., executed on a single thread using instructions that operate on multiple data items.
  • std::execution::par_unseq indicates that a parallel algorithm’s execution may be parallelized, vectorized, or migrated across threads.
In HPX, we have implemented all of the above with correct semantics in the namespace hpx::execution. However, the community quickly realized that the execution policies alone do not provide sufficient flexibility for controlling the (possibly parallel) execution environment. Early discussions in the ISO standardization committee documented the idea of standardizing ‘unified executors’, which was later abandoned [11]. In HPX, we have used the ideas outlined in that paper to develop the extensive customization mechanisms that we describe below.
All of the parallel algorithms in HPX rely on internally invoking a number of customization points that can be overloaded by either the executor or the execution parameters object. We call a customization point a function that can be overloaded by an external user-defined function or object such that it is selected at compile time as a replacement for an internal, predefined default implementation. Customization point objects are usually function object instances that fulfill the two objectives of a) unconditionally trigger (conceptified) type requirements on the arguments of that function, and b) dispatch to the correct function via argument dependent lookup (ADL). In C++, these are often a library feature that adds concept checking resulting in, e.g. clearer compilation error messages in case of erroneous template instantiations. We describe the set of customization points used by HPX’s standard algorithm implementation relevant to this work in Section 4.2. In HPX, we rely on the tag_invoke methodology [12] for the implementation of our customization points.
4.2 Customization Points for HPX’s Parallel Algorithms
As outlined in Section 2, in order to optimize a parallel algorithm, two parameters need to be adaptively controlled: the size of the chunks, i.e. the size of a subsection of the data array that is passed to a single concurrent task, and the number of processing units. Here we assume that the amount of work in the user-supplied loop body is either known or can be measured during the first invocation of the parallel algorithm and this data can be used to compute the optimal parameters.
HPX’s parallel algorithm implementations rely on invoking several customization points. We will focus, in particular, on three of those: measure_iteration, processing_units_count, and get_chunk_size (see [7]). Listing 1 shows a simplified call sequence taken from the algorithm implementation of HPX. All of the customization points are called by passing the execution parameters object params and the executor exec that were bound to the execution policy used for invoking the parallel algorithm (see Section 4.1). The variable loop_body is a function representing the user-supplied loop body, while count holds the overall number of iterations.
// Return the time per iteration for a given
// loop body and overall number of iterations.
auto iteration_duration = measure_iteration(
params, exec, loop_body, count);
// Return number of cores to utilize for a given time
// per iteration and overall number of iterations.
size_t cores = processing_units_count(
params, exec, iteration_duration, count);
// Return the chunk size for the given time per iteration,
// processing unit count, and overall number of iterations
size_t chunk_size = get_chunk_size(params,
exec, iteration_duration, cores, count);
Listing 1: Exemplar invocation sequence of customization points in the implementation of HPX’s parallel algorithms
The expected semantics of the customization points are:
  • measure_iteration should return the time spent execute the loop body.
  • processing_units_count should return the number of processing units the algorithm implementation should utilize.
  • get_chunk_size should return the number of array elements per task.
The default implementations for these customization points splits the work into equally sized chunks while utilizing all available processing units.
In Section 5 we present the results of implementing a customization point object adaptive_core_chunk_size (acc) that exposes the aforementioned customization functions by encapsulating the mathematical model we describe in Section 3. We will show that such an automatic customization is possible without changing the existing APIs implemented by HPX which are fully conforming to the parallel algorithms as defined by the C++ Standard. The required functionality is encapsulated in a simple C++ type that serves as the customization point for HPX’s parallel algorithms, without having to modify the algorithm’s implementation.
5 Experiments and Methodology
Our new executor is adapting the optimal number of cores and chunk size, so we call it adaptive_core_chunk_size or (acc) in this paper. In this execution parameters object, we have incorporated a customization point—measure_iteration—designed to compute the time per loop execution (See Listing 1). The time ($T_{i})$ will be calculated once for each workload, and then will be used to find $T_{1}$ which is the total time for a given number of iteration. HPX runs a benchmark on an empty thread to calculate overhead which is $T_{0}$. Another customization point used in this executor is processing_units_count (listing  1) that uses $T_{1}$ and $T_{0}$ and applies Equation 7 to find the optimal number of cores for each workload. It then uses that value, unless it is more than the maximum available cores in the system, in which case the maximum available cores are used. The last critical customization point in this executor is get_chunk_size which calculates optimal chunk size based on the calculated cores in processing_units_count. This Customization point uses the minimum time per chunk calculated based on Eq 10 and makes sure to have at least that chunk size for the optimal result (see Section 4.2).
For those more familiar with OpenMP than with HPX, the above implementation would provide an automatic way to pick the best value for the num_threads argument to the parallel pragma. Not only will this avoid slowdowns when loops are too small or quick to benefit from parallelism, but it leaves cores available for other parallel tasks should they be needed.
In order to better understand the behavior of map-type algorithms, we conducted a thorough study of the effects of core count and chunk size on the adjacent-difference algorithm, which is memory-bound, and artificial-work which is compute bound.
Initially, we focused on executing the algorithm using executors with different numbers of cores and chunks. The experiments illuminate how different combinations of cores and chunks influence the algorithm’s performance. In the next section, Experiment 1 (see Section 5), we examined the different conditions requiring varying numbers of chunks and cores for a set of workloads. We examined each characteristic (core and chunk size) separately to obtain the best combination. It became evident that for smaller input sizes, optimal performance is achieved with fewer cores and chunks. Conversely, employing more cores and chunks yields better performance for larger input sizes.
The rationale behind this is that for a smaller workload, an increased number of cores and chunks escalates overhead, hindering performance, while for larger workloads with a larger value of $N_{C}$, more chunks enables greater load balancing. To maintain optimal performance, we use an executor that automatically adjusts the optimal number of cores and chunks.
Experiment 1
We evaluate all the performance measurements in this experiment using HPX V1.10.0 on a test machines with Intel Xeon Skylake processors, with 40 cores at 2.4GHz and 96 Gb of main memory, 2 sockets with 20 cores each, with hyperthreading disabled. In these experiments, each processor unit is the same as one core or thread. We used a benchmark to generate data and the result is average of 50 iterations.
In the initial experiment, we kept the number of cores constant while varying the chunk sizes for various workloads. HPX has a very light-weight parallelism with very efficient work stealing [1]. As a result, we expect to see good resource utilization. However, caution is needed, as excessive chunking can introduce significant overhead. The graphs also indicate the best chunking options for each number of cores used. We tested each number of cores with three possible chunking configurations: a number of chunks equal to the number of cores ($C=1$), four times the number of cores ($C=4$), and eight times the number of cores ($C=8$). See Figure 1.
(a) Speedup with 2 processing units (cores)
(b) Speedup with 16 processing units (cores)
(c) Speedup with 32 processing units (cores)
Figure 1: Array size vs. speedup when using different numbers of processing units (cores) for parallelizing the finite-difference algorithm for different numbers of chunks-per-core, C. For comparison, the value of $C$ in these runs behaves like the chunk size argument to OpenMP’s static scheduling algorithm.
In the second experiment, we compared the different fixed core counts using $C=4$ along with the result with our adaptive algorithm using adapted cores and $C=8$ (acc). See Figure 2. This is equivalent to using OpenMP and providing static scheduling and a num_threads argument.
The results show that using all available cores for every workload is not always the best option. The graph indicates that, for smaller workloads, using fewer cores is more effective, while larger workloads benefit from using more cores.
Figure 2: Speedup measured for the adjacent difference algorithm across a range of cores counts and input sizes. We compare executions (for different numbers of cores) with the results measured when using the new adaptive_core_chunk_size (acc) (red line).
Experiment 2
Experiment 2 used 2 different hardwares, one is the same as experiment one, and the other is AMD EPYC processors with 48 cores, 2 sockets with 24 cores each. However, to evaluate the impact of our adaptive_core_chunk_size executor on compute-bound loops, we used an artificial workload instead of adjacent-difference. We then compared the performance of our new executor against the default parallel execution policies (used by both OpenMP and HPX) across varying workloads.
Figure 4, and  4 present this comparison. In these graphs, parallelization begins earlier and scales more rapidly. As before, $T_{1}$ represents the total execution time. However, this algorithm scales differently compared to the previous case (bigger $T_{1}$ for the same input size). As a result, the adaptive_core_chunk_size execution parameters object not only starts parallelization on smaller arrays, but it also uses all available cores in the system earlier than in the default case. As a result, especially very small workloads show significant benefits from dynamic core adjustments with our new executor. Additionally, our executor consistently outperforms other core configurations in larger workloads. This performance gain is attributed to its ability to select optimal chunk sizes compared to the default parallel policies in HPX.
Figure 3: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on an Intel hardware.
Figure 4: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on an AMD hardware.
6 Results
In this section, we present and analyze the findings from our experiments. Our goal was to identify the optimal core-chunk combinations for a map-style algorithm and develop an adaptive_core_chunk_size execution parameters object.
Performance analysis of varying chunk sizes. In the first experiment, we varied chunk sizes while keeping the number of cores constant. The results, illustrated in Figures 1(a)1(b), and 1(c) highlight how performance varies with chunk sizes across various workloads. Having the chunks eight times of number of cores is always better option.
Performance of adaptive chunking. Building on the insights from the two experiment sets, we developed an adaptive_core_chunk_size customization point object. This object dynamically adjusts core and chunk configurations based on workload size. Figure 2 illustrates the performance of various processor counts alongside the results of our executor. The red line (acc) optimally selects the number of cores and chunks for each workload. By dynamically adjusting the number of processors—utilizing fewer for smaller workloads and more for larger ones—this approach demonstrates improved overall performance across all workloads.
It is also worth noting that in Experiment 1, figure 2, although our new executor achieves the best overall performance, the parallel efficiency remains limited, even with 40 cores. Specifically, we observe only approximately a 10x speedup compared to the sequential execution. This is primarily because adjacent_difference is a memory-bound algorithm. However, when the algorithm is changed to a compute-bound one, as shown in figures 4 and 4, the benefits of parallelism become more evident. On a 40-core machine, we achieve up to 38x speedup, and on a 48-core machine, the speedup reaches up to 46x compared to the sequential execution.
In summary, the new approach is beneficial in three aspects: (1) performance gain, the adaptive executor consistently outperformed static configurations. By dynamically tuning core and chunk sizes, it minimized overheads and maximized resource utilization, leading to significant performance improvements across all workload sizes. (2) overhead reduction, for small workloads, the adaptive executor effectively reduced overhead by limiting cores and chunks. For large workloads, it leveraged maximum available cores and increased chunk count, enhancing parallelism. (3) Adaptability, the adaptive executor’s ability to adjust to varying workloads makes it highly versatile, suitable for a wide range of parallel applications.
7 Conclusions
The development and optimization of parallel algorithms are crucial for maximizing computational resource efficiency. This paper studies map-type algorithms and develops a mathematical model to improve their performance based on runtime parameters. The focus of the study is on determining the optimal combination of core counts and chunk sizes at runtime for various workloads. The model has been implemented and integrated in the HPX C++ library. By introducing the HPX execution parameters object adaptive_core_chunk_size, we have made it possible to dynamically adjust the number of cores and the size of the chunks based on the application’s workload size and the available number of cores in the execution environment. The result is improved performance and scalability across a wide range of workloads for parallel algorithms. This adaptive approach reduces the overheads and enhances resource utilization, leading to better overall performance. This is particularly true for smaller workloads.
8 Future Work
Future work will focus on several areas to build upon the findings of this paper. First, we aim to explore the application of the adaptive_core_chunk_size execution parameters object to other algorithms. This includes investigating its impact on additional memory-bound and I/O-bound algorithms to assess its versatility and performance benefits across different computational paradigms. Second, fluctuations in the performance graphs suggest the influence of cache effects. Future models can factor in these effects by leveraging measurement tools like PAPI.
Lastly, we plan to conduct comprehensive benchmarking on a variety of real-world applications to validate the executor’s effectiveness and robustness. This includes performance evaluations on different hardware architectures, such as GPUs and specialized accelerators, ensuring the executor’s adaptability and efficiency across diverse computing environments. By testing on a broader range of platforms, we aim to demonstrate the generalizability and scalability of our approach.